OpenStudy (anonymous):

Please help: Simplify the expression: radical sign 7x times radical sign 14x^3

OpenStudy (anonymous):

\[\sqrt{7x}\times \sqrt{14x^3}\]

OpenStudy (gw2011):

[(7x)^1/2][(14x)^3]=(98x^4)^1/2=(2)^1/2 (49x^4)^1/2=7x^2(2)^1/2